How many yards of material from a 24-yard length of cloth remain after 3 pieces,each 3 1/2 yards long,and 5 pieces,each 2 1/4 long are removed?
How do you solve this?

Answer by boilpoil(127)

24-3(7/2)-5(9/4)
=24-42/4-45/4
=96/4-87/4
=9/4(yard)
(=2 1/4) --- If your teacher or something wants you to write in normal fractions, write this.